a.-b. Merchandise Inventory, before adjustment, has a balance of $6,500. The newly counted inventory balance is $7,000. Unearned Seminar Fees has a balance of $5,000, representing prepayment by customers for five seminars to be conducted in June, July, and August 2019. Two seminars had been conducted by June 30, 2019. Prepaid Insurance has a balance of $6,000 for six months’ insurance paid in advance on May 1, 2019. Store equipment costing $19,840 was purchased on March 31, 2019. It has a salvage value of $400 and a useful life of six years. Employees have earned $150 that has not been paid at June 30, 2019. The employer owes the following taxes on wages not paid at June 30, 2019: SUTA, $4.50; FUTA, $0.90; Medicare, $2.18; and social security, $9.30. Management estimates uncollectible accounts expense at 1 percent of sales. This year’s sales were $1,000,000. Prepaid Rent has a balance of $5,100 for six months’ rent paid in advance on March 1, 2019. The Supplies account in the general ledger has a balance of $300. A count of supplies on hand at June 30, 2019, indicated $100 of supplies remain. The company borrowed $10,600 from First Bank on June 1, 2019, and issued a four-month note. The note bears interest at 6 percent. Required: Based on the information above, record the adjusting journal entries that must be made for Sufen Consulting on June 30, 2019. The company has a June 30 fiscal year-end. Analyze: After all adjusting entries have been journalized and posted, what is the balance of the Prepaid Rent account?
Answer #1
Account Titles Debit Credit
Merchandise Inventory $              500.00 =7000-6500
      Cost of Goods Sold $              500.00
(Increase in inventory on hand)
Unearned Seminar Fees $           2,000.00 =5000/5*2
      Seminar Fees $           2,000.00
(Fees earned during the period)
Insurance Expense $           2,000.00 =6000/6*2
       Prepaid Insurance $           2,000.00
(Prepaid insruance expired)
Depreciation Expense $              810.00 =(19840-400)/6*3/12
      Accumulated Depreciation     $              810.00
(Deprecaition expense for the period)
Wages Expense $              150.00
        Wages Payable $              150.00
(Wages accrued but not paid)
Payroll tax expense $                16.88
       SUTA Payable $                   4.50
       FUTA Payable $                   0.90
       Medicare Payable $                   2.18
       Social Security Payable $                   9.30
(Payroll tax expense)
Bad Debt Expense $        10,000.00 =1000000*1%
      Allowance for Doubtful Debts $         10,000.00
(Bad debt expense)
Rent Expense $           3,400.00 =5100/6*4
        Prepaid Rent $           3,400.00
(Prepaid rent expired during the period)
Supplies Expense $              200.00 =300-100
       Supplies $              200.00
(Supplies consumed during the period)
Interest Expense $                53.00 =10600*6%*1/12
       Interest Payable $                53.00
(Interest accrued but not paid)

Balance in Prepaid Rent = $5100 - 3400 = $1700